There are x teams in a league and each team plays each other team twice, then the total number of games, N, played for the season is: N=x^2-x. If the league played 182 games, how many teams are there?
Answer by checkley71(8403)

N=X~2-X OR 182=X~2-X OR X~2-X-182=0 OR (X-14)(X+13)=0 OR X=14 TEAMS.